OpenStudy (anonymous):

What is the vertex of the function y = 2x2 + 16x + 21? The two after the x is an exponent

OpenStudy (anonymous):

Vertex. Find x by -b/2a. Once you have x plug into equation to get y.

OpenStudy (anonymous):

you need an equation of the form:\[y=a(x-h)^2+k\] where (h,k) is the vertex so complete the square: \[y=2x^2+16x+21\] subtract 21 from both sides:\[y-21=2x^2+16x\] factor the right hand side with 2:\[y-21=2(x^2+8x)\] complete the square by adding 32 to both sides in the parentheses: in this case (16):\[y+11=2(x^2+8x+16)\] now write as perfect square:\[y+11=2(x+4)^2\] now put in vertex form:\[y=2(x+4)^2-11\]

OpenStudy (anonymous):

can u find the vertex using vertex form from there?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

did u get (-4,-11) as your vertex?

OpenStudy (anonymous):

yes ! (:

OpenStudy (anonymous):

Did you get what i did when i completed the square

OpenStudy (anonymous):

because i don't just want u to get the answer without understanding how you got it

OpenStudy (anonymous):

So LOOK at all my steps carefully and take them in

OpenStudy (anonymous):

no i got it we learned this in class i just needed a refresher on how to do it. Thank you for the help !